OpenStudy (anonymous):

who knows what is the equation of directrix of an ellipse? both vertical and horizontal. for (x-h)^2/b^2+(y-k^2/a^2)=1 ???

OpenStudy (anonymous):

fr vertical and hrizntal ellpse pleaaaaaaaaaaaase

ganeshie8 (ganeshie8):

since an ellipse has two focus points, it will have two directrices also

OpenStudy (anonymous):

do you knw what is the equation?

ganeshie8 (ganeshie8):

use that chart

OpenStudy (anonymous):

thank you s much!!!!!!!!!!!!

ganeshie8 (ganeshie8):

np :) that chart has everything you ever need.. when dealing wid ellipses

OpenStudy (anonymous):

thank you very much and have a nice day!

ganeshie8 (ganeshie8):

for vertical ellipses :- two directrix equations : \(\large y = k+ \frac{a^2}{c}\) \(\large y = k- \frac{a^2}{c}\)

ganeshie8 (ganeshie8):

ty :) you too have a great day !!